Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
6x – 4y = 19
24x - 16y = -76

Answer by Alan3354(69443)

6x – 4y = 19
24x - 16y = -76
----------------
6x – 4y = 19
6x - 4y = -19 2nd eqn times 1/4
---------------- Subtract
0 = 38
--> inconsistent, no solution